Grade 12Mechanics

A rigid body is acted upon by a horizontal force which is inversely proportional to the dist covered s . the work done by this force will b proportional to??

Bhavna Kochrekar
9 Years agoGrade 12

3 Answers

Khimraj
ApprovedApproved Tutor Answer9 Years ago
Work done by force is give by W = ∫ F.dsAccording to question Force F =k/sWhere k is constant so work done isW = ∫ (k/s)ds = klns So work is is directly proportional to lns
Adarsh Kumar
4 Years ago

F=sk

work done is given by dw=F.ds
=>dw=sk.ds
integrating both sides we get w=k lns.
so, work is directly proportional to k lns
 
 
Adarsh Kumar
4 Years ago
F= K/S
​ work done is given by dw=F.ds
=>dw= k/s × ds
integrating both sides we get w=k lns.
so, work is directly proportional to klns